OpenStudy (jjuden):

Which is the direct linear variation equation for the relationship? y varies directly with x and y = 12 when x = 4.

OpenStudy (jjuden):

y = 2x + 4 y = 3x y = x + 8 y = x – 8

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

you got ps3

OpenStudy (jjuden):

no i have xbox 360 way better

OpenStudy (anonymous):

2

OpenStudy (anonymous):

or B

OpenStudy (jjuden):

thank you some to help me with math not some annoying loser

OpenStudy (anonymous):

no probs

OpenStudy (anonymous):

oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes oh yes

OpenStudy (jjuden):

omg little kid go away

OpenStudy (anonymous):



OpenStudy (anonymous):

hey by the way im 13 and my abs look like this

OpenStudy (anonymous):